Astronomy & Astrophysics General Information

Springer Verlag

Dear author,

The Board of Directors has decided that from now on all manuscripts, including letters, submitted to A&A have to be written in L^{A}TEX with Springer’s L^{A}TEX 2ε A&A macro package.

You can obtain the A&A macro package via our anonymous ftp server

host: ftp.springer.de (192.129.24.12),

  directory: /pub/tex/latex/aa

You can also send an email to svserv@vax.ntp.springer.de containing the line:

  send /tex/latex/aa.zip

For advice on TEX problems please send an email to texhelp@springer.de.

Authors are required to submit two hardcopies for the referees (this requires the referee option to be set in the \documentclass declaration) to either of the editors:

J. Lequeux Astronomy and Astrophysics Editorial Office Observatoire de Paris 61 avenue de l’Observatoire F-75014 Paris - France Tel. 33-(0)1 43 29 05 41 Fax 33-(0)1 43 29 05 57

email: aanda@mesiob.obspm.fr

H.J. Habing

Astronomy and Astrophysics

Editorial Office Leiden Observatory P.O. Box 9513 2300 RA Leiden - The Netherlands Tel. (31) (0) 71-5 27 59 16 Fax (31) (0) 71-5 27 58 03

                              email: aanda@strw.leidenuniv.nl

Letter Editor: P.Schneider Max-Planck-Institute für Astrophysik Karl-Schwarzschild-Strasse 1 D-85740 Garching - Germany Tel. (49)(0)(89) 32 99 32 96 Fax (49)(0)(89) 32 99 35 96

email: aal@mpa-garching.mpg.de

After the paper has been accepted, send the final manuscript (remove the referee option) as *.tex (with the figures as *.ps or *.eps files and eventually additional stylefiles) via ftp (ftp.springer.de; z230; password: spvz230) or email to astro@springer.de or via conventional mail to:

Springer Verlag Journal Production Physics / 3141 Ulrike Hadek / Petra Meyer Tiergartenstr. 17 D-69121 Heidelberg – Germany Fax: (0)6221 487 180

Please note that in the future A&A will appear in its as usual printed form and also as a parallel-to-print edition on the Internet:

  http://www.springer.de/phys/journals/journals.html